OpenStudy (anonymous):

Quilt squares are cut on the diagonal to form triangular quit pieces. The hypotenuse of the resulting triangles is 34 inches long. What is the side length of each piece?

OpenStudy (anonymous):

@whpalmer4 can you help me ?

OpenStudy (whpalmer4):

Okay, this problem is the exact opposite of the problem we just did. If you have a square and cut it diagonally, the hypotenuse is exactly \(\sqrt{2}\) times the length of the sides of the square.

OpenStudy (whpalmer4):

If the length of the side is \(s\), \[34 = s\sqrt{2}\]Can you solve for the value of \(s\)?

OpenStudy (anonymous):

im confused

OpenStudy (whpalmer4):

Okay, tell me what is confusing you...

OpenStudy (whpalmer4):

 \[s^2 + s^2 = 34^2\]

OpenStudy (anonymous):

the whole 34 = 8 \[\sqrt{2}\]

OpenStudy (whpalmer4):

 \[1^2 + 1^2 = c^2\]\[1+1=c^2\]\[2=c^2\]\[\sqrt{2} = c\] So, in such a triangle (diagonal of a square), the hypotenuse is always \(\sqrt{2}\) times the length of the side. We know the hypotenuse of this triangle is 34. We therefore know that \[s\sqrt{2} = 34\]where \(s\) is the length of a side of the square. \[s\sqrt{2} = 34\]\[\frac{s\sqrt{2}}{\sqrt{2}} = \frac{34}{\sqrt{2}}\]\[s = \frac{34}{\sqrt{2}}\]

OpenStudy (anonymous):

its not 34 radical of 2 is it ?

OpenStudy (whpalmer4):

How can the legs of a right triangle be longer than the hypotenuse?!? The hypotenuse is the longest side of the triangle. \(\sqrt{2}\approx 1.414\) so anything you multiply by it gets larger, not smaller.

OpenStudy (anonymous):

would it be 17 radical 2 ?

OpenStudy (whpalmer4):

Do you know how to rationalize radical expressions?

OpenStudy (whpalmer4):

\[\frac{34}{\sqrt{2}}\]is a radical expression with a radical in the denominator. When we rationalize it, we remove the radical from the denominator by multiplying both numerator and denominator by the denominator. That's just a fraction that equals 1, so it doesn't change the value at all: \[\frac{34}{\sqrt{2}} * \frac{\sqrt{2}}{\sqrt{2}} = \frac{34\sqrt{2}}{\sqrt{2}*\sqrt{2}}\]Can you simplify that any more?

OpenStudy (anonymous):

yes ?

OpenStudy (whpalmer4):

Okay, please do so!

OpenStudy (anonymous):

I cant remember how to simplify that

OpenStudy (whpalmer4):

what is \(\sqrt{2}*\sqrt{2}=\)

OpenStudy (anonymous):

1 ?

OpenStudy (whpalmer4):

do you remember what I said the value of \(\sqrt{2}=\)

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

No. Scroll back and read.

OpenStudy (anonymous):

1.414

OpenStudy (whpalmer4):

in case you don't know, \(\approx\) means "approximately" — the value of the square root of 2 is a never-ending string of digits, so I can't write \(\sqrt{2} = 1.414...\) because I couldn't write them all down.

OpenStudy (whpalmer4):

Right. So what is \(1.414*1.414=\)

OpenStudy (anonymous):

1.999396

OpenStudy (anonymous):

you there ?

OpenStudy (whpalmer4):

yes, what number is that very close to?

OpenStudy (whpalmer4):

\[\sqrt{2}*\sqrt{2} = 2\]The definition of the square root of 2 is that number which when multiplied by itself gives you 2. \[\frac{34*\sqrt{2}}{\sqrt{2}*\sqrt{2}} = \frac{34*\sqrt{2}}{2} = 17\sqrt{2}\]

OpenStudy (whpalmer4):

You seem very shaky on this material — do you have a teacher you can ask for some help? If not, the Khan Academy videos have been very good, in my experience. https://www.khanacademy.org/math/algebra/exponent-equations/simplifying-radical-expressions/v/radical-equivalent-to-rational-exponents-2